What is the balanced chemical reaction for (NH4)2SnCl6 and H2O. then the reaction with heat. What is the new balanced chemical reaction if (NH4)2SnCl6 is replaced by (NH4)2PbCl6?
Answer #1

1

(NH4)2Sn Cl6 +H2O = (NH4)2SnCl4 +2HCl+ O-

2

(NH4)2Pb Cl6 + H2O = PbCl4+ (NH4)2O +2HCl

(NH4)2PbCl6 +H2O =H2PbCl6   - ( NH4)2O =PbCl4 +2 HCl

(I hope this will be correct)